OpenStudy (anonymous):

Jake is now ¼ of his father’s age and in 7 years time his age will be 1/3 rd of his father. How old is Jake now?

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

1/4x

OpenStudy (anonymous):

can i show u how i did my equation and then u can correct me?

OpenStudy (anonymous):

(X+4x)=4(x+4) Idk why I did that

OpenStudy (anonymous):

im kinda confused i dont know what i did

OpenStudy (anonymous):

i dont know where to add the 7

Parth (parthkohli):

If Jake's father is \(\rm f\), then Jake is \(\rm {j \over 4}\). Agree?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Okay, and after 7 years, Jake's father will be \(\rm f + 7\) right?

Parth (parthkohli):

What I meant above was that Jake is \(\rm f \over 4\)

OpenStudy (anonymous):

yeah

Parth (parthkohli):

Also, Jake will be \(\rm {f \over 4} + 7\) right?

OpenStudy (anonymous):

yeah

Parth (parthkohli):

Now you have that in 7 years, Jake will be 1/3rd of his father. That means that if you multiply Jake's age in 7 years by 3, you get his father's age in 7 years.

OpenStudy (anonymous):

oh ok that makes sense

Parth (parthkohli):

Making an equation for that,\[\rm 3\left({f \over 4} +7\right) = f + 7\]Get how I made this equation?

OpenStudy (anonymous):

yes i get it now thanx alot :)

Parth (parthkohli):

You're welcome! P.S: This question is actually not linear algebra... it's just algebra!

OpenStudy (anonymous):

oh really lol i didnt know that

Parth (parthkohli):

Heh. I used to get confused too when I was a basic dude.

OpenStudy (anonymous):

man that makes me feel better haha

Parth (parthkohli):

;0